OpenStudy (anonymous):

H12P1: CURRENT SOURCE if A=100 v=5V, i=9mA ±VS=17V 500Ω within 0.1% of the exact. I got: a)555.555555556 b)1333.33333333 c) ??????? What is ans of c=????

OpenStudy (nurali):

0.008833240684

OpenStudy (nurali):

if correct please click best response.

OpenStudy (anonymous):

how to calculate R for part b?

OpenStudy (anonymous):

I=Vout/R1+R2 R1=R2

OpenStudy (anonymous):

nurali...plzzzz help Suppose that we need to put precisely i=9mA through a load that presents a resistance but the value of the resistance is not well controlled. Suppose also that we have a precision voltage source of v=5V, but this voltage source is pretty wimpy: it cannot supply more than 0.1mA without smoke coming out of it. With the circuit shown above we can choose a resistor of resistance R that can accomplish this goal. First, assuming that the op amp is ideal (A=∞) give the value of the resistance R, in Ohms, that we need.

OpenStudy (anonymous):

for part 1 R1=v/i

OpenStudy (nurali):

tell me value Vs.

OpenStudy (anonymous):

i have the values Vs=17V

OpenStudy (anonymous):

hey for part c..... answer will be from the R of part a i=(Vi/R(1+((R+Rl)/R*A))

OpenStudy (anonymous):

tell me the value of Rl so that i could find values for your question's part C